# Enhancing Visual Reasoning with Autonomous Imagination in Multimodal Large Language Models

### v1

@@ 3.2 Imagination Space
To handle both 2D and 3D visual inputs, we developed separate 2D and 3D imagination spaces, where 2D space uses image as representation and 3D space is based on the 3D Gaussian splatting model [19]. The imagination space is designed to support visual rendering and supports a minimal set of operators: focus, ignore, and transform. Focus isolates relevant content for further manipulation, ignore enables MLLMs to disregard extraneous information, and transform allows the repositioning of desired content. This compact set of operations enables MLLMs to reason and solve practical challenges by themselves, as demonstrated in our experiments. MLLMs select operators through natural language output, which are then applied to the imagination space. The updated space is rendered and returned to MLLMs for the subsequent reasoning step.
@@ 22             $\mathcal{G}_{\text{marked}}\leftarrow\mathcal{G}_{\text{marked}}\cup\{g_{i}\}$
3.3 Focus Operator
The focus operator allows MLLMs to automatically isolate and label the target content from a scene, creating distinct elements for further transformation. Once MLLMs identify an object of interest using a virtual cursor (as described in Sec. 3.6), the focus operator segments this content for focused manipulation.
In 2D space, this cursor directly conditions the Segment Anything Model (SAM) [20] to perform segmentation. MLLMs verify segmented output to ensure alignment with their intended focus. However, in 3D Gaussian space, segmentation on demand poses unique challenges: Existing methods [62, 46] are designed for unconditional segmentation, where all objects are segmented without the ability to specify conditions. However, our use case requires conditional segmentation, focusing on a specific object selected by an input condition, making these methods unsuitable.
To enable the focus operator in 3D Gaussian-based representations, we introduce a method to selectively segment an object from an unstructured 3D Gaussian scene, which is illustrated in Alg. 1. Given the MLLM’s initial selection block, we first generate a virtual camera trajectory that orbits around the targeted object, creating a video sequence. This sequence is fed into the SAM2 model [42], which applies the selection cursor condition across all rendered frames. For each frame, rays are cast from pixels within the segmentation mask, and path tracing records the radiance contributions of intersecting Gaussians. This process is expressed as volumetric integration illustrated in Line 4-11. Gaussians that receive a contribution higher than a threshold will receive one vote.
We select 3D Gaussian with votes exceeding 10% of the highest vote, forming a shell around the target object’s radiance field. This shell, along with the internal radiance Gaussians, is segmented collectively to define the object’s radiance structure. In both 2D and 3D spaces, the segmented object is placed in a separate layer, while the remaining content is retained in a base layer.
@@ 3.4 Ignore Operator
The ignore operation removes a focused object from the imagination space to prevent interference with the reasoning process. Since API access to state-of-the-art MLLMs doesn’t support attention masks, removing the object leaves a hole in the base layer, potentially causing hallucinations. To address this, we simply inpaint the hole created by the object’s removal.
For both 2D and 3D imagination spaces, we project the object mask onto the rendered image as the region to inpaint. We directly use the implementation provided by OpenCV [15] by simply merging the nearby pixels. The inpainted image is then used for further reasoning by the MLLMs, effectively avoiding it from paying attention to the inpainted region as the content has been removed.
@@ 3.5 Transform Operator
Object transformation can occur in four directions: up, down, left, and right, each represented by an alphabet character $(a,b,c,d)$ to avoid interference from semantically meaningful words. In each step of the iterative imagination process, MLLMs select a direction for movement. Since MLLMs struggle with determining precise distances, we standardize movement units in screen space coordinates and gradually reduce step size throughout the process, where MLLMs only control the direction of movement.
@@ 3.6 Reasoning Process
Specifically, the iterative reasoning process described in Sec. 3.1 is executed as follows: After initializing the imagination space, a virtual cursor is positioned at the center and automatically considered as focused. MLLMs can perform transform operations on the cursor to reposition it within the space. Following each operation, MLLMs conduct scene modification and receive an updated image of the newly rendered imagination space.
@@
When MLLMs identify that the cursor has selected a visual element of interest, a focus operation is performed based on the cursor’s location. If MLLMs determine that this operation has correctly segmented the intended element, the focus is shifted to that object, enabling MLLMs to execute either a transformation or ignore operation on it. Once MLLMs have either repositioned or disregarded the object, the focus returns to the virtual cursor. This process iterates continuously as part of the reasoning workflow until all necessary reasoning steps are completed, culminating in a final textual response.
4 Benchmark and Evaluation Protocol
While previous works have demonstrated MLLMs’ capabilities in various visual tasks, a significant gap remains in their ability to perform practical, intuitive reasoning tasks that are simple for humans. Rather than benchmarking existing abilities, our focus is on expanding MLLMs’ capabilities, enabling them to tackle tasks that were previously unsolvable. To this end, we propose the Intuitive Visual Reasoning Benchmark (IVRB), which evaluates three tasks that require practical visual reasoning: dense object counting, simple jigsaw puzzle solving, and object placement.
@@ 4 Benchmark and Evaluation Protocol
Dense Object Counting: When faced with numerous densely packed objects, humans often rely on multi-step reasoning to reach an accurate count, as a single glance may not suffice. Existing MLLMs, however, struggle with dense object counting, as their multi-step reasoning is not yet effective for this task. We include this task to assess whether the model can leverage reasoning to offset its limited direct perception, mirroring human strategies.
@@
For evaluation, we directly compare the model’s predicted count to the ground truth. In addition to reporting the success rate, we calculate the mean and variance of counting errors to provide insight into each model’s accuracy and consistency. We constructed in total 122 images with paired ground truth for our benchmark.
@@
Solving Simple Jigsaw Puzzles: Jigsaw puzzles are classic tests of visual perception and reasoning, commonly used to assess intelligence. In this task, we evaluate MLLMs’ ability to solve simple jigsaw puzzles, aligning their problem-solving performance with that of humans to assess visual reasoning capabilities. In the Jigsaw Puzzle Solving task, MLLMs are tasked with identifying and placing missing pieces in their correct locations.
@@
For evaluation, we use a digital jigsaw puzzle game with a magnetic mechanism, where pieces automatically snap into place when positioned close to their correct locations. Success is measured by the completion rate, defined as the percentage of pieces placed in their correct locations. For evaluation, we constructed 11 cases where four pieces are missing, and 11 cases where six pieces are missing. The jigsaw puzzle’s size ranges from $3\times 5$ to $5\times 8$.
@@
Object Placement: Previous methods have shown progress in enabling MLLMs to understand and describe static scenes, such as identifying object locations (e.g., "Where is the cup?"). However, for practical use, MLLMs must also interpret dynamic instructions that convey intent, such as "Where should the cup be placed?" In the Object Placement task, MLLMs are required to identify both the current and target locations of objects based on abstract instructions (e.g., "Prepare two cups for guests in the living room"). Given the input scene of 3D Gaussians, MLLMs must determine the original locations of objects and their intended locations based on a provided prompt.
@@
5 Experiments
@@ 5.1 Baselines
We evaluate our approach against state-of-the-art MLLM, namely GPT-4o [38]. We provide a clearly defined 2D coordinate system to GPT-4o to ensure output coordinates are generated without any ambiguity. We also adopt the recently proposed representative clue finding method VCoT [45] using their open-released pre-trained model. We developed a baseline GPT-4o Sampling inspired by the sample-then-evaluate imagination paradigm used in robotics [18, 8]. Since their original methods are designed for robotic manipulation, we re-implement them under our paradigm. Note that this sampling strategy is not designed for general visual reasoning, and can only be applied to some of the tasks in our evaluation. Furthermore, to demonstrate the necessity of altering the visual scene beyond simply finding visual clues, we crafted a robust reasoning baseline named cursor-only that utilizes our imagination space but restricts operations solely to transform operations of the virtual cursor. Note that this baseline serves as an ablation of our method, which has the same closed-loop control design except that scene modification is disabled and only virtual cursor moving is enabled.
5.2 Dense Counting
As shown in Tab. 1, our model significantly outperforms all baselines, achieving a higher success rate and lower errors when mistakes occur. Note that the sampling method is not directly applicable to the counting task due to intractability so we omit its comparisons. VCoT only supports drawing one visual marker for clue finding, hence is unsuitable for more complex reasoning tasks involving multiple objects, such as counting. We also compare with our cursor-only baseline in counting, observing that it performs similarly to GPT-4o. We further experimented with reasoning by drawing multiple visual markers, implemented by allowing cursor-only baseline to draw visual markers when performing the focus operation. Although technically able to draw markers on every ball to perform counting accurately, the model quickly enters a negative feedback loop: hallucinations lead it to draw more markers, which introduce additional noise and exacerbate the hallucinations. This results in a success rate of 0% and intractable mean error and variance, as the model cannot stop counting, further highlighting the importance of modifying the visual scene.
We further conduct analysis to increase the counting difficulty by adding more objects with partial data. Results are shown in Fig. 5(a) with an additional baseline GPT-4o-text-cot, which performs counting via pure text-based CoT. This reveals perceptual limitations in GPT-4o that cannot be overcome by simply increasing the textual CoT steps, while our method remains unaffected. Fig. 5(b) shows the percentage of cases solved as reasoning steps increase. For reference, we include the one-step results from GPT-4o depicted as a flat line. Our method performs consistently better as reasoning steps grow, whereas GPT-4o-text-cot fails to do this and may even introduce noise, resulting in poorer performance than one-step GPT-4o.
Settings GPT-4o VCoT Ours (cursor-only) GPT-4o Sampling Ours
Dense Counting Success Rate 39.8% 15.5% 39.0%/0% - 85.3%
Mean Error 0.82 2.90 1.02/intractable - 0.17
Variance 1.49 15.57 2.54/intractable - 0.22
Simple Jigsaw Puzzle 4-Piece Missing 29.5% 9.1% 27.3% 43.2% 68.2%
6-Piece Missing 9.1% 3.3% 24.2% 30.3% 51.5%
Object Placement Locating 10.9% 10.4% 69.4% - 69.4%
Placement 3.6% 1.5% 27.8% 17.3% 37.3%
@@ 5.2 Dense Counting
Table 1: Quantitative comparison results. See Sec. 4 for details of the metrics. We consider two variants of cursor-only in dense counting, leading to two sets of results. Furthermore, GPT-4o Sampling is intractable in some tasks as illustrated. See Sec. 5.2 and Sec. 5.4 for details. In object placement, cursor-only functions the same as our method in locating, so their results are identical.
@@
(c) Counting by highlighting visual clues only
Figure 5: We demonstrate in (a)(b) that as counting task difficulty increases linearly, scaling inference-time textual reasoning (implemented as GPT-4o-text-cot, See Sec. 5.2) fails—and even performs worse than vanilla GPT-4o—as complexity exceeds perception limits. In contrast, our methods remain unaffected by these limits, achieving correct counting even as difficulty rises. Additionally, highlighting visual clues can introduce noise, causing MLLMs to enter hallucination loops; for instance, in (c), the model incorrectly concluded there were 196 balls when only two were present. Qualitative comparisons of counting are provided in the appendix.
@@ 5.3 Simple Jigsaw Puzzle Solving
As shown in Tab. 1 and Fig. 6, our method consistently outperforms all baselines. It achieves a higher locate accuracy, which contributes to a higher success rate with the same number of attempts. Since this visual reasoning task requires the model to choose the correct target location rather than merely locating visible elements, VCoT is ineffective at reasoning. The sampling method performs reasonably well in this task since we set a large sampling budget to ensure the correct answers can be sampled as the candidate answers. However, despite its extensive reasoning budget, it still does not match the performance of our method. This highlights the effectiveness and robustness of establishing an autonomous perception-control loop. We also demonstrate that the cursor-only baseline performs similarly to primitive GPT-4o under the four-piece missing scenario. However, it performs significantly better under the more challenging six-piece missing scenario, further confirming the effectiveness of the closed-loop control paradigm.

### v3

@@ 3.3 Imagination Space
The imagination space is designed to render images for MLLMs to perceive and supports a minimal set of operators for MLLMs to call: focus, ignore, and transform. Focus isolates relevant content for further manipulation, ignore enables MLLMs to disregard extraneous information, and transform allows the repositioning of desired content. This compact set of operations enables MLLMs to reason and solve practical challenges by themselves, as demonstrated in our experiments. MLLMs select operators through natural language output, which are then applied to the imagination space. The updated space is rendered and returned to MLLMs for the subsequent reasoning step.
We discuss the detailed implementation of the operators in the following sections. The operators are implemented separately for the 2D space (represented as images) and the 3D space (represented via 3D Gaussians). MLLMs are unaware of this implementation difference, as it interacts exclusively with rendered images as input and executes operators identically in both scenarios.
3.4 Focus Operator
The focus operator allows MLLMs to isolate and label target content from a scene, creating distinct elements for further transformation. Once MLLMs identify an object of interest using a virtual cursor (as described in Sec. 3.7), the focus operator segments this content for focused manipulation. In 2D space, this cursor directly conditions the Segment Anything Model (SAM) (Kirillov et al., 2023) to perform segmentation. MLLMs verify segmented output to ensure alignment with their intended focus. In the 3D Gaussian space, segmentation on demand poses unique challenges: Existing methods (Ye et al., 2024; Shen et al., 2024) are designed for unconditional segmentation, where all objects are segmented without the ability to specify conditions. However, our use case requires conditional segmentation, focusing on a specific object selected by an input condition, making these methods unsuitable.
To enable the focus operator in 3D Gaussian-based representations, we introduce a method to selectively segment an object from an unstructured 3D Gaussian scene, which is illustrated in Alg. 1. Given the MLLM’s initial selection block, we first generate a virtual camera trajectory that orbits around the targeted object, creating a video sequence. This sequence is fed into the SAM2 model (Ravi et al., 2024), which applies the selection cursor condition on all rendered frames. For each frame, rays are cast from pixels within the segmentation mask, and path tracing records the radiance contributions of intersecting Gaussians. This process is expressed as volumetric integration illustrated in Line 4-11. Gaussians who receive a contribution higher than a threshold will receive one vote.
We select the 3D Gaussian with votes exceeding 10% of the highest vote, forming a shell around the target object’s radiance field. This shell, along with the internal radiance Gaussians, is segmented collectively to define the object’s radiance structure. In both 2D and 3D spaces, the segmented object is placed in a separate layer, while the remaining content is retained in a base layer.
@@ 3.5 Ignore Operator
The ignore operation removes a focused object from the imagination space to prevent interference with the reasoning process. Since API access to state-of-the-art MLLMs does not support attention masks, removing the object leaves a hole in the base layer, potentially causing hallucinations. To address this, we simply inpaint the hole created by the removal of the object. For both 2D and 3D imagination spaces, we project the object mask onto the rendered image as the region to inpaint. We directly use the implementation provided by OpenCV (Itseez, 2015) by simply merging the nearby pixels. The inpainted image is then used for further reasoning by the MLLMs, avoiding paying attention to the inpainted region as the content has been removed.
@@ 3.6 Transform Operator
Object transformation can occur in four directions: up, down, left, and right, each represented by an alphabet character $(a,b,c,d)$ to avoid interference with semantically meaningful words. In each step of the iterative imagination process, MLLMs select a direction for movement. Since MLLMs struggle with determining precise distances, we standardize movement units in screen space coordinates and gradually reduce step size throughout the process, where MLLMs only control the direction of movement.
@@ 3.7 Reasoning Process
The iterative reasoning process described in Sec. 3.2 is executed as follows: After initializing the imagination space, a virtual cursor is positioned at the center and automatically considered as focused. MLLMs can perform transform operations on the cursor to reposition it within the space. Following each operation, MLLMs conduct scene modification and receive an updated image of the newly rendered imagination space.
@@
When MLLMs identify that the cursor has selected a visual element of interest, a focus operation is performed based on the location of the cursor. If MLLMs determine that this operation has correctly segmented the intended element, the focus is shifted to that object, enabling MLLMs to execute either a transformation or ignore operation on it. Once MLLMs have either repositioned or disregarded the object, the focus returns to the virtual cursor. This process iterates continuously as part of the reasoning workflow until all necessary reasoning steps are completed, culminating in a final textual response.
4 Experiments
In the experiments, we challenge MLLMs with four tasks that are intuitive to humans and require strong visual understanding and reasoning abilities: counting, simple jigsaw puzzle solving, object placement, and multi-object hallucination. We target at verifying that compared to 1) simply scaling textual reasoning steps 2) conducting visual prompting, while still conducting visual-to-textual conversion within a single step, the closed-loop task decomposition paradigm is indeed more effective and necessary.
4.1 Benchmark and Evaluation Protocol
@@
Counting: When faced with numerous densely packed objects, humans often rely on multi-step reasoning to reach an accurate count, as a single glance may not suffice. Existing MLLMs, however, struggle with counting, as their multistep reasoning is not yet effective for this task. We include this task to assess whether the model can leverage reasoning to offset its limited direct perception, mirroring human strategies.
@@
For evaluation, we directly compare the predicted count of the model with the ground truth. In addition to reporting the success rate, we calculate the mean and variance of the counting errors to provide insight into the accuracy and consistency of each approach. We constructed 122 images with real objects with paired ground truth.
@@
Solving Simple Jigsaw Puzzles: Jigsaw puzzles are classic tests of visual perception and reasoning, commonly used to assess intelligence. In this task, we evaluate the ability of MLLMs to solve simple jigsaw puzzles, aligning their problem-solving performance with that of humans to assess visual reasoning capabilities. In the jigsaw puzzle solving task, MLLMs are tasked with identifying and placing missing pieces in their correct locations.
@@
For evaluation, we use a digital jigsaw puzzle game with a magnetic mechanism, where pieces automatically snap into place when positioned close to their correct locations. Success is measured by the completion rate, defined as the percentage of pieces placed in their correct locations. For evaluation, we constructed 11 cases where four pieces are missing and 11 cases where six pieces are missing. The size of the puzzle ranges from $3\times 5$ to $5\times 8$.
@@
Object Placement: Previous methods have shown progress in enabling MLLMs to understand and describe static scenes, such as identifying object locations (e.g., “Where is the cup?”). However, for practical use, MLLMs must also interpret dynamic instructions that convey intent, such as “Where should the cup be placed?” In the object placement task, MLLMs are required to identify both the current and target locations of objects based on abstract instructions (e.g., “Prepare two cups for guests in the living room”). Given the input scene of 3D Gaussians, MLLMs must determine the original locations of objects and their intended locations based on a provided prompt.
@@
Multi-Object Hallucination: Current MLLMs can suffer from hallucination by perceiving or generating non-existing objects in the input scene. The complexity of the input scene is beyond the limitation in perceptual capability of MLLMs, which is a plausible reason for hallucination. Our approach can serve as a promising way to address this challenge, especially for the scenes where multiple objects exist. By the closed-loop iterative process of autonomous imagination, MLLMs can deal with objects one by one, as in our counting benchmark, instead of handling them all together, effectively reducing the possibility of hallucination. We conducted additional experiments on the recently proposed multi-object hallucination benchmark ROPE (Chen et al., 2024) to validate this argument. We focus on the most challenging subset in the benchmark where the state-of-the-art model on the official leaderboard, GPT-4o, exhibits the poorest performance, which is answering about single object given multiobject image, heterogeneous object types, on unseen data split. We refer to the detailed benchmark setting from the original paper.
@@ 4.2 Baselines
We evaluate our approach against state-of-the-art MLLM, namely GPT-4o (OpenAI, 2024a) (our approach is also utilized to enhance GPT-4o in all experiments). We provide a clearly defined 2D coordinate system to GPT-4o so that output coordinates are generated without any ambiguity. We also adopt the recently proposed visual reasoning method VCoT (Shao et al., 2024) using their open-released pre-trained model. We developed a baseline GPT-4o Sampling inspired by the sample-then-evaluate imagination paradigm used in robotics (Kapelyukh et al., 2024; Ding et al., 2024). Since their original methods are designed for robotic manipulation, we reimplement them under our method. Note that this sampling strategy can only be applied to some of the tasks in our evaluation since sampling in large solution spaces is heavily intractable. Furthermore, to demonstrate the necessity of altering the visual scene for task decomposition beyond adding visual prompts, we crafted a reasoning baseline named cursor-only that utilizes our imagination space, but restricts operations solely to transform operations of the virtual cursor. Note that this baseline serves as an ablation of our method, which has the same closed-loop control design except that scene modification is disabled and only virtual cursor moving is enabled. Furthermore, we also adopt two recently proposed advanced MLLMs, namely OpenAI’s o1 (OpenAI, 2024b) and Molmo (Deitke et al., 2024). o1 is well known to have specifically trained reasoning abilities under pure textual modality. We utilize it to justify whether simply scaling text-time textual reasoning could solve the tasks. On the other hand, Molmo is trained to have ability to utilize visual markers as cues in visual modality for reasoning. We utilize it to justify the necessity of task decomposition instead of conducting visual prompting with visual markers.
@@
Table 1: Quantitative comparison results under counting, jigsaw puzzle solving, and object placement. See Sec. 4.1 for details of the metrics. We consider two variants of cursor-only in counting, leading to two sets of results, and GPT-4o Sampling cannot be applied to counting and locating as illustrated, please see Sec. 4.3, 4.2, and Sec. 4.5 for details. In object placement, cursor-only functions the same as our method in locating, so their results are identical. Furthermore, NA* indicates that under these tasks, Molmo constantly fails to output reasonable coordinates or refuses to output.
Settings GPT-4o o1 Molmo VCoT Ours (cursor-only) GPT-4o Sampling GPT4o+Ours
Counting Success Rate 39.8% 50.8% 86.2% 15.5% 39.0%/0% - 85.3%
Mean Error 0.73 0.74 0.34 3.11 1.12/not calculable - 0.19
Variance 0.62 0.90 2.57 5.51 1.22/not calculable - 0.22
Simple Jigsaw Puzzle 4-Piece Missing 29.5% 31.8% NA* 9.1% 27.3% 43.2% 68.2%
6-Piece Missing 9.1% 27.3% NA* 3.3% 24.2% 30.3% 51.5%
Object Placement Locating 10.9% 17.3% NA* 10.4% 69.4% - 69.4%
Placement 3.6% 8.5% NA* 1.5% 27.8% 17.3% 37.3%
4.3 Counting
The results are shown in Tab. 1. Our approach significantly improves the performance of original GPT-4o, achieving a higher success rate and lower errors when mistakes occur. It is worth noting that Molmo is known to be specifically trained to conduct counting with its native “pointing” ability. Our approach enables GPT-4o to perform on par with it. Note that the sampling method is not directly applicable to the counting task due to intractability, so we omit its comparisons. VCoT only supports drawing one bounding box to highlight a piece of visual evidence, hence is unsuitable for more complex reasoning tasks involving multiple objects, such as counting. We also compare with our cursor-only baseline in counting, observing that it performs similarly to GPT-4o. Since having a cursor moving around might be ineffective for counting, we adopt a cursor-only variant, where the MLLM can draw a bounding box on the cursor location to highlight the object during counting. Although technically able to draw boxes on every ball to perform counting accurately, the model quickly enters a negative feedback loop: hallucinations lead it to draw more markers, which introduce additional noise and exacerbate the hallucinations. This results in a success rate of 0% and makes mean error and variance not calculable, as the model cannot stop counting, further highlighting the importance of modifying the visual scene.
To further verify the essentialness of task decomposition in visual modality, we conduct additional analysis under the real object data by increasing the counting difficulty by adding more objects with partial data. The purpose was to evaluate the impact of difficulty and reasoning steps on performance. To achieve this, we selected a subset of cases from the whole data and supplemented them with simpler cases involving fewer objects. We then re-conducted the experiments on this partial dataset using GPT-4o, alongside a newly introduced textual CoT reasoning baseline. In the first experiment (Fig. 4(a)), we show the success rate of different methods when faced with varying numbers of objects. In the second experiment (Fig. 4(b)), we test how different methods improve when given different limits on the number of reasoning steps. The result is shown as the solvable rate, which indicates the percentage of cases the method can solve under the current budget. When imposing such a limit, our model is restricted to performing no more than the specified number of steps, with each step defined as a combination of the focus operation and the subsequent operation performed after focusing. For the textual CoT baseline, we explicitly instructed the model on the maximum number of reasoning steps it was allowed to use. The results are shown in Fig. 4(a) with an additional baseline GPT-4o-text-cot, indicating the pure text-based CoT. This reveals perceptual limitations in GPT-4o that cannot be overcome by simply increasing the textual CoT steps, while our method remains unaffected. Fig. 4(b) shows the percentage of cases solved as reasoning steps increase. For reference, we include the one-step results from GPT-4o depicted as a flat line. Our method performs consistently better as reasoning steps grow, whereas GPT-4o-text-cot fails to do this and may even introduce noise, resulting in poorer performance than one-step GPT-4o.
@@ 4.3 Counting
(c) Counting by drawing bounding boxes only
Figure 4: (a)(b) show that as counting task difficulty increases linearly, scaling inference-time textual reasoning (implemented as GPT-4o-text-cot, See Sec. 4.3) fails—and even performs worse than vanilla GPT-4o—as complexity exceeds perception limits. In contrast, our methods remain unaffected by these limits, achieving correct counting even as difficulty rises. Additionally, adding visual prompts can introduce noise, causing MLLMs to enter hallucination loops; for instance, in (c), the model incorrectly concluded there were 196 balls when only two were present. The qualitative results of counting are provided in the appendix.
@@ 4.4 Simple Jigsaw Puzzle Solving
In the jigsaw puzzle solving task, the rotations of pieces are not considered. As multiple trials are allowed, we restrict the total number of attempts to 20 and report the finish rate, defined as the proportion of missing pieces successfully placed. The sampling method involves comparing pairs of target locations and iteratively selecting the better option until a single target location remains. We guarantee that the correct answer exists within the sampling process. In our method, once a jigsaw piece is moved to the imagined target location, it is considered to be placed. GPT-4o and o1 select the target location by outputting coordinates, while VCoT and Molmo determine the target location by drawing a block or marker, leveraging their specialized training for such tasks.
As shown in Tab. 1 and Fig. 5, our method consistently outperforms all baselines. It achieves a higher locating accuracy, which contributes to a higher success rate with the same number of attempts. Since this visual reasoning task requires the model to choose the correct target location rather than simply locating visible elements, VCoT is ineffective at reasoning. The sampling method performs reasonably well in this task, since we set a large sampling budget to ensure the correct answers can be sampled as the candidate answers. However, despite its extensive reasoning budget, it still does not match the performance of our method. This highlights the effectiveness and robustness of establishing an autonomous perception-control loop. We also demonstrate that the cursor-only baseline performs similarly to primitive GPT-4o in the four-piece missing scenario. However, it performs significantly better under the more challenging six-piece missing scenario, further confirming the effectiveness of our closed-loop control approach.